The rational function f( 2/-2×-5)



A .Find f¹(×)



B. Find the equation of the line of symmetry of f¹(×) and f(×).


1
Expert's answer
2022-05-16T16:12:44-0400

A.


"f(x)=\dfrac{2}{-2x-5}"

"-2x-5\not=0=>x\not=-2.5"

"y=\dfrac{2}{-2x-5}"

Interchange "x" and "y"


"x=\dfrac{2}{-2y-5}"

Solve for "y"


"2y+5=-\dfrac{2}{x}"

"y=-\dfrac{1}{x}-\dfrac{5}{2}"

"y=-\dfrac{2+5x}{2x}"

Replace "y" with "f^{-1}(x)"


"f^{-1}(x)=\dfrac{-2-5x}{2x}"



B.

The line of symmetry of "f^{-1}(x)" and "f(x)" is "y=x."